Use point slope form to find equation of line, then write in in slope intercept form
Y-10=1/5 (x-1)

Answer by stanbon(75887)

Use point slope form to find equation of line, then write in in slope intercept form
Y-10=1/5 (x-1)
-----
y = (1/5)x - (1/5) + 10
-----
y = (1/5)x + (49/5)
------------
Cheers,
Stan H.
------------